Operation Summary for September 10, 2010 - September 13, 2010 |
The Integrated Luminosity for the period from 9/6/10 to 9/13/10 was 48.41 pb-1. Friday Sept 10 The day shift began with Operations monitoring the Tevatron (TeV) store 8077 (202.74E30) and the Recycler stash (7.6E10), with the Antiproton Source (Pbar) stacking (55.9E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, RF personnel accessed Booster to work on a RF station (BRF5) and they accessed into MI to replace the power amplifier on MIRF18. Kicker personnel accessed MI to check fluorinert levels in the kickers and to work on the MI-10 injection kicker. The MI fluorinert checks were completed by 12:45 PM. The Booster RF work was completed by 1 PM. Operations resumed sending beam to MiniBooNE. Kicker personnel completed their work on the MI injection kicker at 2:45 PM, MI resumed mixed mode stacking at 3 PM. During the evening shift, the NuMI autotune program had some issues; operators started the tomcat servlets, which fixed this problem. FESS completed repairs to the Booster East Gallery chilled water return line. Saturday Sept 11 The midnight shift began with Operations monitoring store 8077 (51.55E30) and the stash (249.3E10), with Pbar stacking (8.3E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8077 at 3:52 AM. ACL script problems had delayed the end of the store. Operations put the TeV into HEP shot setup and established store 8078 at 5:32 AM with an average initial luminosity (AIL) of 248.4E30. ACL script problems continued through shot setup; Operations consulted with TeV specialists about workarounds. During the day shift, operators monitored systems. During the evening shift, Operations terminated store 8078 at 5:30 PM, put the TeV into shot setup, and established store 8079 at 6:46 PM with an AIL of 245E30. CDF had problems turning on that delayed their initial luminosity report by ~30 minutes. Operators tuned the Proton Source. Sunday Sept 12 The midnight shift began with Operations monitoring store 8079 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, at 6:32 AM, store 8079 aborted just before its scheduled termination time. The beam aborted at TeV sector B3 due to vacuum issues. Investigation into the cause lasted until just before the end of the shift, when operators put the TeV through a dry squeeze. Operators tuned Linac and MI to improve efficiencies. During the day shift, operators called in vacuum controls personnel to look at the B3 CIA crate; they ended up changing the sector valve control card, which had erroneously pulled the abort. The 400 MeV line started tripping; operators traced this to sparking in the 400 MeV chopper power supply. Operators called in support personnel to make repairs. Operations put the TeV into shot setup at 10:01 AM and established store 8081 at 11 AM with an AIL of 231.49E30. Operators search and secured the Meson enclosures MO1 through M05 and turned on power supplies. During the evening shift, Operations terminated store 8081 at 10:33 PM, put the TeV into shot setup, and established store 8082 at 11:41 PM with an AIL of 252.7E30. Monday Sept 13 The midnight shift began with Operations monitoring store 8082 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations turned off the MiniBooNE beam at 5 AM in preparation for an access during the day shift. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 09/14/2010 email Fermilab |
| Security, Privacy, Legal | |